OpenStudy (giovxnnx):

Mean or Proportion?! 1. Chips Ahoy wants to perform a study to determine the number of chocolate chips in their cookies. To that end, they collected a sample of 40 cookies. The mean of this sample is 23.95 chocolate chips. From past studies, we know that the standard deviation is 2.55 chocolate chips. Construct a 99% confidence interval of the mean of chocolate chips in all such cookies. The situation refers to a mean or proportion? 2. As a manager for an advertising company, you must plan a campaign designed to increase Twitter usage. A recent survey suggests that 85% of adults know what Twitter is. How many adults should you survey in order to be 90% confident that your estimate is within 5% of the true population proportion? The situation refers to a mean or proportion?

OpenStudy (perl):

Here is the formula for confidence interval of a mean\[ \Large \overline x \pm z_{\alpha/2}\frac{\sigma}{\sqrt n}\]

OpenStudy (perl):

Do you know what the symbols represent. Let me know, I will explain.

OpenStudy (mathmale):

The quantity\[z _{\frac{ \alpha }{ 2 }} \] is called the "z-critical value." Know how to find this value for a 99% confidence interval?

OpenStudy (mathmale):

If not, would you mind looking up "z critical value" on the Internet or in your textbook?

OpenStudy (giovxnnx):

I am so lost, I still cannot understand. :(

OpenStudy (agent0smith):

Perl gave the formula. mathmale explained the z critical value. \(\large \overline x \) is the mean. \(\large \sigma\) is the standard deviation \(\large n\) is the sample size. Read the question, identify each value, then plug them into the formula.